It is known that 60% of households in South Africa own at least one dog. What is the mean number of dogs per household in South Africa?


It is known that 60% of households in South Africa own at least one dog. A survey is carried out on 20 households. The probability that 12 households visited have at least one dog is?


1
Expert's answer
2022-03-27T16:00:43-0400

"EX\geqslant 0.6\cdot 1=0.6-only\,\,estimate\,\,can\,\,be\,\,made, because\,\,there\,\,is\\no\,\,information\,\,about\,\,the\,\,percent\,\,of\,\,people\,\,having\,\,2,3,... dogs\\\\The\,\,number\,\,of\,\,households\,\,having\,\,at\,\,least\,\,one\,\,dog\,\,Y\sim Bin\left( 20,0.6 \right) \\P\left( Y=12 \right) =C_{20}^{12}\cdot 0.6^{12}\cdot 0.4^8=0.179706"